# Gravitational waves from strong first order phase transitions

### v1

@@ V.2 Comparison to simulations
In Fig. 5 (top row) we present the shear stress power spectra at a single time, along with those obtained from the convolutions of velocity power spectra (see Eqs. 47, 48). We convolve with 3-velocity spectra and weighted 4-velocity spectra. The sum of the convolution contributions is close in shape to the measured spectra, but the weighted 4-velocity spectra are closer to the measured amplitude, particularly for the detonation. This motivates the use of the weighted 4-velocity in the gravitational wave production calculation. The remaining mismatch signals a departure from Gaussianity for the weighted 4-velocity field.
Figure 5: Top row: comparison of the measured shear stress power spectrum normalised by energy density squared ${\mathcal{P}}_{\Pi}/\bar{\epsilon}^{2}$ with that obtained by convolving the measured weighted 4-velocity power spectra (see Eq. 47, 48; in dashed lines) and the 3-velocity power spectra (in dash-dotted lines). Left panel: detonation ($\alpha_{\text{n}}=0.67$, $v_{\text{w}}=0.92$) at time $3600/T_{c}$; right panel: deflagration ($\alpha_{\text{n}}=0.5$, $v_{\text{w}}=0.44$) at time $4400/T_{c}$. Bottom row: comparison of the instantaneous rate of change of the gravitational wave power spectra $\dot{\mathcal{P}}_{\text{gw}}$ computed via the convolution model prediction (68). Left panel: detonation ($\alpha_{\text{n}}=0.67$, $v_{\text{w}}=0.92$) with growth between times $2400/T_{c}$ and $3600/T_{c}$ ; right panel: deflagration ($\alpha_{\text{n}}=0.5$, $v_{\text{w}}=0.44$) with growth between times $3600/T_{c}$ and $4400/T_{c}$. Convolution predictions are shown with dashed lines in four colours, representing the three terms (purely compressional, mixed compressional-vortical, and purely vortical) and their sum. In the bottom row we show with a dotted line the sound shell model prediction, including the phenomenological suppression factor [32], whose value is given in the legend.
@@
$\bar{R}=R_{*}\left[{\xi_{U_{\parallel}}(t_{1})\xi_{U_{\parallel}}(t_{2})}/\xi^{2}_{U_{\parallel}}(t_{*})\right]^{1/2}$ (76)
@@
We see that both the amplitude and shape are good agreement with the convolution prediction, with perhaps better agreement than the shear stress power spectrum. We might expect to see the same differences between convolution prediction and simulation as in the shear stress power spectrum, as the GW spectrum is obtained by a linear operation from the shear stress UETC. There are several approximations made to go from the shear stress UETC to the GW spectrum, so the better agreement is perhaps a fortuitous cancellation of the non-Gaussianity with errors in the approximations.
@@
Figure 6: Extrapolations of power spectra of gravitational waves over time. Left-hand-side panels correspond to the detonation ($\alpha_{n}=0.67$, $v_{w}=0.92$) the right-hand side ones to the deflagration ($\alpha_{n}=0.5$, $v_{w}=0.44$).Top panels show fits to Eq. (77) (faded orange) to ${\mathcal{P}}_{\text{gw}}(k,t)$ (solid orange) over time for selected $kR_{*}$ in the interval $kR_{*}\in[8.6,80.9]$. The fitted growth index $d(k)$ in the function (77) for each gravitational wave power at each individual $k$-mode in the range $kR_{*}\in[8,80]$ is shown in the panels on the second row. In the third row is the extrapolated late time power ${\mathcal{P}}_{\text{gw}}(k)$ (dashed black, grey time varying uncertainties) along with the power spectra at the last available time in the (orange). In the bottom panels are shown the total (solid black) and partial (solid orange) gravitational wave power, fits (dotted lines) to the function in Eq. (78), and the extrapolated total power (dashed black, with grey fit uncertainty band).
In the same figure we also show the sound shell model prediction of each transition computed via the PTtools package [76] in its default precision, with the suppression factor of Ref. [32] included. In the case of the detonation, we see the model does not accurately predict the growth rate of the gravitational wave power spectrum, signaling a breakdown in assumptions behind the model. The deflagration however, shows very good agreement (after the correction by the suppression factor) between sound shell model and simulation for $kR_{*}\leq 10^{2}$.
To further test the sound shell model, we compared its growth rate prediction for the weak phase transitions studied in Ref. [30] ($\alpha_{\text{n}}=0.0046$, and the same wall speeds) with the numerical simulations, shown in Fig. 10 in Appendix E. The agreement at wavenumbers near the peak and higher is very good, both in amplitude and shape. At lower wavenumbers the simulations do not have enough volume or duration to be safely compared.
@@
In Fig. 6 (middle row) we show the extrapolation of the fits as $t\to\infty$ (that is, the parameter ${\mathcal{P}}_{\text{gw}}^{\infty}$), along with a fit uncertainty estimated from the parameter covariance matrix. For low wavenumbers, where there are few modes in the wavenumber bin, the oscillations in the gravitational wave power are large and the estimate more uncertain. At high wavenumber ($kR_{*}\gg 1$), the asymptote is well determined, and the power spectrum is close to convergence.
Phase transition $d_{\text{tot}}$ $12\zeta+2\lambda-$ 1
@@
In Fig. 6 (bottom row) we show the growth of the partial gravitational wave power along with the prediction obtained by integrating the fitted form of the power spectrum (77).
@@
The theoretical expectation is that the approach indexes $12\zeta+2\lambda-$ 1 should be in agreement with the fitted $d_{\text{tot}}$. The agreement is reasonably good for the deflagration, but very poor for the detonation. However, the power law approach indices of individual modes are in much better agreement with $12\zeta+2\lambda-$ 1 (see Fig. 6). We note that the total power is dominated by modes which are close to the box size in wavelength, and the power law indices of volume-averaged quantities are likely to suffer from significant finite-volume effects.
@@
$\alpha_{\text{n}}=0.5$, $v_{\text{w}}=0.44$ $0.017\pm 0.003$
@@
The similar magnitudes of the dimensionless constants $\tilde{\Omega}_{\text{gw}}^{\infty}$ suggests that there is an approximate efficiency factor of gravitational wave production of order $10^{-2}$ throughout the lifetime of the flow. A factor with a similar order of magnitude was observed in the early growth rates of gravitational wave power for transitions of weaker strength ($\alpha_{\text{n}}=0.0046,0.05$) in Ref. [30]. This supports the idea that, in the case $H_{\text{n}}t_{\text{sh}}\ll 1$ studied here, a reasonable estimate of the gravitational wave power can be obtained by multiplying the initial growth rate by the shock decay time $t_{\text{sh}}=R_{*}/\bar{U}_{\parallel,*}$ [10]. A similar efficiency factor, although about a factor 2 larger, was obtained at $\alpha_{\text{n}}=0.5$ in Ref. [35]. Reasons for the difference include: smaller simulations ($512^{3}$), a different definition of shock decay time (called the source duration $\tau_{\text{sw}}$ there), and a different definition of the reference kinetic energy fraction. The simulations were also carried out for walls speeds $v_{\text{w}}\leq 0.8$ at which the bubble solutions are deflagrations or hybrids. For these solutions, the Higgsless method misses the wall deceleration effect when compression waves start to interact, and overestimates the kinetic energy fraction. Nonetheless, one might expect to see the same efficiency factor if one normalises with the observed kinetic energy.
@@ VI Conclusions
We studied two cases, a detonation with wall speed $v_{\text{w}}=0.92$ and transition strength parameter of $\alpha_{\text{n}}=0.67$, and a deflagration characterised by $v_{\text{w}}=0.44$ and $\alpha_{\text{n}}=0.5$. These two cases provide contrasting examples of phase transitions, as they have different proportions of compressional and vortical modes in the total fluid velocity, as was seen in [32]. Here, we evolved on much larger grids, and for a sufficiently long time after the phase transitions to allow the non-linearities to affect the evolution.
We studied the evolution of the velocity power spectra, separating compressional and vortical components, and the shear stress spectrum, the source of the gravitational wave power spectrum. We also measured separately, for the first time, the spatial components of the 4-velocity weighted by the enthalpy density. The shear stress is quadratic in these components. Another first is the study of the unequal time correlations of the velocity field and the shear stress, allowing us to study the time decorrelation functions. The decorrelation functions are an important part of the modelling of the gravitational wave power spectrum, which as the key observable, we also followed.
In the detonation, non-linearities become clear even before collisions of the sound shells: the enthalpy-weighted RMS 4-velocity is approximately 40% higher than the RMS 3-velocity, and there is a substantial change in the weighted 4-velocity profile after the collision. The collisions generate vorticity, but the RMS velocity in the vortical modes is roughly an order of magnitude smaller than that of the compressional modes.
In the deflagration, the RMS velocity is smaller, and below the naive value estimated from the sound shells, as observed in [32]. The RMS 3-velocity and RMS weighted 4-velocity are similar, and there is more equal division between compressional and vortical modes, which appears to evolve towards a constant ratio, around $1.4$ in this case, with the compressional mode dominating. Turbulence seems to be developing here, although as the simulation does not run for much more than one eddy turn-over time, we do not see the emergence of the $k^{-2/3}$ Kolmogorov spectrum, observed in freely decaying vortical turbulence [31]. Instead, the vortical velocity spectrum shows a broad and approximately flat ($k^{0}$) plateau, which is still quite different from the characteristic $k^{-1}$ of the compressional modes. The larger length scale in the plateau is the mean bubble spacing, and we speculate that the smaller length scale is the thickness of the vortices produced by the collision of the sound shells.
@@
The envelope of the compressional mode oscillations is well fitted by a Gaussian in the wave number multiplied by the time difference, with a coefficient $V_{\parallel}$ with dimensions of velocity. In the deflagration, this parameter is around $0.06$, which is close to $v_{\perp}/\sqrt{3}$, the RMS of the vector components of the vortical flow velocity. This would be the expectation in a model where the compressional modes are “swept” [48] by the vortical modes. On the other had, in the detonation, the coefficient $V_{\parallel}$ is twice $v_{\perp}/\sqrt{3}$. It is also about half $v_{\parallel}/\sqrt{3}$, so the origin of the Gaussian envelope is unclear. One possibility is through interactions of shocks themselves, which are non-trivial (see e.g. [70, 71, 72]). We see $\lambda$-configurations, three-way junctions of shocks, are clearly visible in the slices through the simulation volume of the detonation (Fig. 1).
@@
The shear stress power spectrum has quite different shapes between the detonation and deflagration, as a consequence of the importance of the vortical modes in the latter. The shapes are very close to those obtained by a convolution of the weighted 4-velocity spectra, except that the overall amplitude is higher This points to significant non-Gaussianity in the flow, particularly for the detonation. There, the convolution with the weighted 4-velocity gives a significant improvement in the agreement over convolution with the 3-velocity. The decorrelation time of the shear stress at wavenumber $k$ is controlled by the dominant fluid oscillation mode at that wavenumber: around $\pi/c_{\text{s}}k$ for flow which is mostly compressional at that wavenumber, and around $1/k\bar{U}$ for mostly vortical. The shear stress in the deflagration is mostly vortical in a wavenumber range $20\lesssim kR_{*}\lesssim 200$.
@@
Quantitatively, we computed the gravitational wave power spectrum with the assumption of a Gaussian velocity field. The computation convolves products of two velocity power spectra with kernels derived from integrating the gravitational wave Green’s function with the decorrelation functions. There are three possible contributions: pure compressional, mixed compressional-vortical, and pure vortical. The result is a prediction for the growth rate of the gravitational wave power spectrum $\Delta\mathcal{P}_{\text{gw}}/\Delta t$ between the two times at which the velocity field is evaluated. The result in both the detonation and the deflagration is that both the amplitude and the shape of the growth rate are well predicted by the compressional modes alone. This is perhaps a surprise given the importance of the vortical modes in the shear stress spectrum. We draw the conclusion that vortical motion does not seem to source a significant contribution to gravitational wave production in first order phase transitions of strength $\alpha_{\text{n}}\lesssim 1$.
We can understand the relative efficiency of acoustic production as a resonance effect: there is a range of sound wave wavenumber pairs $({\mathbf{q}},\tilde{{\mathbf{q}}})$ for which the shear stress oscillates at the same frequency $c_{\text{s}}(q+\tilde{q})$ as the resulting gravitational wave, $|{\mathbf{q}}+\tilde{{\mathbf{q}}}|$. The shear stress for a vortical mode of a given wavenumber has a spread of frequencies, of order $\overline{V}_{\!\perp}q$, which are well below resonance for the gravitational wave.
@@
Our work has important implications for theoretical modelling, for example the sound shell model of [37, 36]. While considering only the compressional modes appears to be a good approximation, the assumption that sound shells do not interact is not correct. This is already known for deflagrations, where the interaction leads to a slowing of the phase boundary, and the eventual formation of shrinking hot droplets, which do not leave much kinetic energy behind [32, 60]. Here, a phenomenological suppression factor, which needs further investigation, gives a reasonable approximation.
@@
Theoretical modelling also needs to account for the kinetic energy decay, and further non-linear shape evolution, rather than assuming a stationary flow. Kinetic energy decay has recently been studied numerically [35], and power laws close to the predicted $t^{-10/7}$ [34] have been observed. There is also predicted to be an associated $t^{2/7}$ evolution of the integral scale, which was not included in Ref. [35]. Other simulations [34] show that the compressional mode spectrum evolves towards a self-similar form, which can be fitted with a broken power law with asymptotic indices $k^{5}$ at low wave numbers and $k^{-1}$ at high wave numbers. The evolution towards this shape also needs to be investigated, as the majority of the gravitational wave emission will probably occur during the evolutionary phase.
@@
Our results for the decorrelation functions also have implications for future modelling. In current implementations, it is assumed that the decorrelation functions of compressional modes oscillate indefinitely with period set by sound speed, leading to a kernel which (in the limit $kR_{*}\to\infty$) can be approximated by a $\delta$-function. The new form given in Eq. 64 improves the prediction of the rate of gravitational wave power spectra. The ratio between a $M_{\parallel}=1.14$ and $M_{\parallel}$ for the $C[U_{\parallel},U_{\parallel}]$ is contained in the range $[0.4,2.0]$, where the lowest ratios occur for the lowest values of $kR_{*}$ (below order 10), while the largest closer to $kR^{*}\approx 9\times 10^{2}$.
@@
$\Omega_{\text{gw},0}/(H_{\text{n}}R_{*})^{2}=$
=$=\left\{\begin{array}[]{cc}(4.8\pm 1.1)\times 10^{-8}&(v_{\text{w}}=0.92,\alpha_{\text{n}}=0.67)\\ (1.3\pm 0.2)\times 10^{-8}&(v_{\text{w}}=0.44,\alpha_{\text{n}}=0.50)\\ \end{array}\right.$ (81)
@@
We conclude with a remark on the deflagration. As previously mentioned, due to a reheating effect that slows down the rate of the phase transition greatly, enacting a reduction of wall speed in tandem $v_{\text{w}}\approx 0.4\rightarrow v_{\text{w}}\approx 0.1$. In effect, and although this reduces the observed gravitational wave power as was explained in [32], it may also provide a way to enact baryogenesis in the latter part of the phase transition. Electroweak baryogenesis [79] in its original form was argued to be most efficient in transitions with low asymptotic wall speeds, at odds with strong gravitational wave production. However, this reheating implies that electroweak baryogenesis and gravitational wave production need not be mutually exclusive: the two different regimes (prior and after reheating) could allow for a sufficiently strong (to be observed) gravitational wave imprint, while still allowing for baryogenesis to occur. For example, the deflagration we simulate produces a loud signal for sufficiently large $H_{\text{n}}R_{*}$, while approximately 30% of the universe is swept by walls with speed $v_{\text{w}}\approx 0.1$, which if maintained from the beginning would produce a much weaker signal. A detailed exploration of this effect as function of transition strength, wall speed and bubble size deserves further investigation.

### v2

@@ V.2 Comparison to simulations
In Fig. 5 (top row) we present the shear stress power spectra at a single time, along with those obtained from the convolutions of velocity power spectra, Eqs. (47, 48). We convolve with 3-velocity spectra and weighted 4-velocity spectra. The sum of the convolution contributions is close in shape to the measured spectra, but the weighted 4-velocity spectra are closer to the measured amplitude, particularly for the detonation. This motivates the use of the weighted 4-velocity in the gravitational wave production calculation. The remaining mismatch signals a departure from Gaussianity for the weighted 4-velocity field.
Figure 5: Top row: comparison of the measured shear stress power spectrum (solid lines) normalised by energy density squared ${\mathcal{P}}_{\Pi}/\bar{\epsilon}^{2}$ with that obtained by using Eqs. (47, 48) to convolve the measured weighted 4-velocity power spectra (dashed lines) and the 3-velocity power spectra (dash-dotted lines). Left panel: detonation ($\alpha_{\text{n}}=0.67$, $v_{\text{w}}=0.92$) at time $3600/T_{c}$; right panel: deflagration ($\alpha_{\text{n}}=0.5$, $v_{\text{w}}=0.44$) at time $4400/T_{c}$. Bottom row: comparison of the measured instantaneous rate of change of the gravitational wave power spectra $\dot{\mathcal{P}}_{\text{gw}}$ (solid lines) with that computed via the convolution model prediction, Eq. (69) (dashed lines). Left panel: detonation ($\alpha_{\text{n}}=0.67$, $v_{\text{w}}=0.92$) with growth between times $2400/T_{c}$ and $3600/T_{c}$ ; right panel: deflagration ($\alpha_{\text{n}}=0.5$, $v_{\text{w}}=0.44$) with growth between times $3600/T_{c}$ and $4400/T_{c}$. For both the top and bottom rows, the convolution predictions are shown with dashed lines in four colours, representing the three terms (purely compressional, mixed compressional-vortical, and purely vortical) and their sum. In the bottom row we show with a dotted line the sound shell model prediction, including the phenomenological suppression factor [37], whose value is given in the legend.
@@
$\bar{R}=R_{*}\left[{\xi_{U_{\parallel}}(t_{1})\xi_{U_{\parallel}}(t_{2})}/\xi^{2}_{U_{\parallel}}(t_{*})\right]^{1/2}.$ (77)
@@
We see that both the amplitude and shape are in good agreement with the convolution prediction, indeed perhaps better agreement than the shear stress power spectrum. We might expect to see the same differences between convolution prediction and simulation as in the shear stress power spectrum, as the GW spectrum is obtained by a linear operation from the shear stress UETC. There are several approximations made to go from the shear stress UETC to the GW spectrum, so the better agreement is perhaps a fortuitous cancellation of the non-Gaussianity with errors in the approximations.
@@
Figure 6: Extrapolations of power spectra of gravitational waves over time. Left-hand-side panels correspond to the detonation ($\alpha_{n}=0.67$, $v_{w}=0.92$); the right-hand side ones to the deflagration ($\alpha_{n}=0.5$, $v_{w}=0.44$). Top panels show fits of ${\mathcal{P}}_{\text{gw}}(k,t)$ (solid orange) to Eq. (78) (faded orange) over time for selected $kR_{*}$ in the interval $kR_{*}\in[8.6,80.9]$. The fitted growth index $d(k)$ in Eq. (78) for each gravitational wave power at each individual $k$-mode in the range $kR_{*}\in[8,80]$ is shown in the panels on the second row. In the third row is the extrapolated late-time power ${\mathcal{P}}_{\text{gw}}(k)$ (dashed black, with fitting uncertainties added) along with the power spectra at the last available time in the simulation (orange). In the bottom panels are shown the total (solid black) and partial (solid orange) gravitational wave power, fits (dotted lines) to the function in Eq. (79), and the extrapolated total power (dashed black, with grey uncertainty band).
In the same figure we also show the sound shell model prediction of each transition computed via the PTtools package (see Ref. [50]) at its default precision level, with the suppression factor of Ref. [37] included. In the case of the detonation, we see the model does not accurately predict the growth rate of the gravitational wave power spectrum, signaling a breakdown in assumptions behind the model. The deflagration however, shows very good agreement between the sound shell model and the simulation for $kR_{*}\leq 10^{2}$, provided the phenomenological supression factor from Ref. [37] is applied. This result validates the supression factor, although we note that we do not have new physical insight into its origin (its relationship with droplets was investigated in Ref. [38]).
To further test the sound shell model, we compared its growth rate prediction for the weak phase transitions studied in Ref. [53] ($\alpha_{\text{n}}=0.0046$, and the same wall speeds) with the numerical simulations, shown in Fig. 10 in Appendix E. The agreement at wavenumbers around the peak and above is very good, both in amplitude and shape. At lower wavenumbers the simulations do not have enough volume or duration to be safely compared.
@@
In Fig. 6 (middle row) we show the extrapolation of the fits as $t\to\infty$ (that is, the parameter ${\mathcal{P}}_{\text{gw}}^{\infty}$), along with a fit uncertainty estimated from the parameter covariance matrix. For low wavenumbers, where there are fewer modes in the wavenumber bin, the oscillations in the gravitational wave power are larger and the estimate more uncertain. At higher wavenumbers ($kR_{*}\gg 1$), the asymptote is well determined, and the power spectrum is close to convergence.
Phase transition $d_{\text{tot}}$ $2\zeta+2\lambda-1$
@@
In Fig. 6 (bottom row) we show the growth of the partial gravitational wave power $\Omega_{par}(t)$ (obtained from integrating data over the restricted wavemode range $kR_{*}\in[8.6,80.9]$) along with the prediction obtained by integrating the fitted form of the power spectrum, Eq. (78).
@@
The theoretical expectation is that the approach indexes $2\zeta+2\lambda-1$ should be in agreement with the fitted $d_{\text{tot}}$. The agreement is reasonably good for the deflagration, but very poor for the detonation. However, the power law approach indices of individual modes are in much better agreement with $2\zeta+2\lambda-1$ (see Fig. 6). We note that the total power is dominated by modes which are close to the box size in wavelength, and the power law indices of volume-averaged quantities are likely to suffer from significant finite-volume effects.
@@
$\alpha_{\text{n}}=0.50$, $v_{\text{w}}=0.44$ $0.017\pm 0.003$
@@
The similar magnitudes of the dimensionless constants $\tilde{\Omega}_{\text{gw}}^{\infty}$ suggests that there is an approximate efficiency factor of gravitational wave production of order $10^{-2}$ throughout the lifetime of the flow. A factor with a similar order of magnitude was observed in the early growth rates of gravitational wave power for transitions of weaker strength ($\alpha_{\text{n}}=0.0046,\,0.05$) in Ref. [53]. This supports the idea that, in the case $H_{\text{n}}t_{\text{sh}}\ll 1$ studied here, a reasonable estimate of the gravitational wave power can be obtained by multiplying the initial growth rate by the shock decay time $t_{\text{sh}}=R_{*}/\bar{U}_{\parallel,*}$ [26]. A similar efficiency factor, although about a factor 2 larger, was obtained at $\alpha_{\text{n}}=0.5$ in Ref. [25]. Reasons for the difference include: smaller simulations ($512^{3}$), a different definition of shock decay time (called the source duration $\tau_{\text{sw}}$ there), and a different definition of the reference kinetic energy fraction. The simulations were also carried out for wall speeds $v_{\text{w}}\leq 0.8$ at which the bubble solutions are deflagrations or hybrids. For these solutions, the Higgsless method misses the wall deceleration effect when compression waves start to interact, and overestimates the kinetic energy fraction. Nonetheless, one might expect to see the same efficiency factor if one normalises with the observed kinetic energy.
@@ VI Conclusions
We studied two cases, a detonation with wall speed $v_{\text{w}}=0.92$ and transition strength parameter of $\alpha_{\text{n}}=0.67$, and a deflagration characterised by $v_{\text{w}}=0.44$ and $\alpha_{\text{n}}=0.5$. These two cases provide contrasting examples of phase transitions, as they have different proportions of compressional and vortical modes in the total fluid velocity, as was seen in Ref. [37]. Here, we evolved on much larger grids, and for a sufficiently long time after the phase transitions to allow the non-linearities to affect the evolution.
We investigated the evolution of the velocity power spectra—separating compressional and vortical components—and the shear stress spectrum, the source of the gravitational wave power spectrum. We also measured separately, for the first time, the spatial components of the 4-velocity weighted by the enthalpy density. The shear stress is quadratic in these components. Another first is the study of the unequal time correlations of the velocity field and the shear stress, allowing us to study the time decorrelation functions. The decorrelation functions are an important part of the modelling of the gravitational wave power spectrum, which as the key observable, we also followed.
In the detonation, non-linearities become clear even before collisions of the sound shells: the enthalpy-weighted RMS 4-velocity is approximately 40% higher than the RMS 3-velocity. The collisions generate vorticity, but the RMS velocity in the vortical modes is roughly an order of magnitude smaller than that of the compressional modes.
In the deflagration, the RMS velocity is smaller, and below the naive value estimated from the sound shells, as observed in Ref. [37]. The RMS 3-velocity and RMS weighted 4-velocity are similar, and there is more equal division between compressional and vortical modes, which appears to evolve towards a constant ratio, around $1.4$ in this case, with the compressional mode dominating. Turbulence seems to be developing here, although as the simulation does not run for much more than one eddy turn-over time, we do not see the emergence of the $k^{-2/3}$ Kolmogorov spectrum, observed in freely decaying vortical turbulence [13]. Instead, the vortical velocity spectrum shows a broad and approximately flat ($k^{0}$) plateau, which is still quite different from the characteristic $k^{-1}$ of the compressional modes. The larger length scale in the plateau is the mean bubble spacing, and we speculate that the smaller length scale is the thickness of the vortices produced by the collision of the sound shells.
@@
The envelope of the compressional mode oscillations is well fitted by a Gaussian in the wave number multiplied by the time difference, with a coefficient $V_{\parallel}$ with dimensions of velocity. In the deflagration, this parameter is around $0.06$, which is close to $v_{\perp}/\sqrt{3}$, the RMS of the vector components of the vortical flow velocity. This would be the expectation in a model where the compressional modes are “swept” by the vortical modes [46]. On the other had, in the detonation, the coefficient $V_{\parallel}$ is twice $v_{\perp}/\sqrt{3}$. It is also about half $v_{\parallel}/\sqrt{3}$, so the origin of the Gaussian envelope is unclear. One possibility is through interactions of shocks themselves, which are non-trivial (see e.g. Refs. [31, 41, 47]). We see $\lambda$-configurations, three-way junctions of shocks, are clearly visible in the slices through the simulation volume of the detonation (Fig. 1).
@@
The shear stress power spectrum has quite different shapes between the detonation and deflagration, as a consequence of the importance of the vortical modes in the latter. The shapes are very close to those obtained by a convolution of the weighted 4-velocity spectra, except that the overall amplitude is higher This points to significant non-Gaussianity in the flow, particularly for the detonation. There, the convolution with the weighted 4-velocity gives a significant improvement in the agreement over convolution with the 3-velocity. The decorrelation time of the shear stress at wavenumber $k$ is controlled by the dominant fluid oscillation mode at that wavenumber: around $\pi/c_{\text{s}}k$ for flow which is mostly compressional at that wavenumber, and around $1/k\bar{U}$ for mostly vortical. The shear stress in the deflagration is mostly vortical in the wavenumber range $20\lesssim kR_{*}\lesssim 200$.
@@
To compare with the gravitational wave power spectrum obtained directly from simulations, in Section V we computed the gravitational wave power spectrum with the assumption of a Gaussian velocity field. The computation convolves products of two velocity power spectra with kernels derived from integrating the gravitational wave Green’s function with the decorrelation functions. There are three possible contributions: pure compressional, mixed compressional-vortical, and pure vortical. The result is a prediction for the growth rate of the gravitational wave power spectrum $\Delta\mathcal{P}_{\text{gw}}/\Delta t$ between the two times at which the velocity field is evaluated. The result in both the detonation and the deflagration is that both the amplitude and the shape of the growth rate are well predicted by the compressional modes alone. This is perhaps a surprise given the importance of the vortical modes in the shear stress spectrum. We draw the conclusion that vortical motion does not seem to source a significant contribution to gravitational wave production in first order phase transitions of strength $\alpha_{\text{n}}\lesssim 1$.
[…]
@@
Our work has important implications for theoretical modelling, for example the sound shell model of Refs. [49, 55]. While considering only the compressional modes appears to be a good approximation, the assumption that sound shells do not interact is not correct. This is already known for deflagrations, where the interaction leads to a slowing of the phase boundary, and the eventual formation of shrinking hot droplets, which do not leave much kinetic energy behind [37, 38]. Here, a phenomenological suppression factor (from Ref. [37]), which needs further investigation, gives a reasonable approximation. The presence of non-linearities in the detonation post-collision is also suggested by the disappearance of the ”knee” at $kR_{*}\geq 10^{2}$ present in the Sound Shell Model prediction 5, a scale associated with the thickness of the rarefaction wave.
@@
Theoretical modelling also needs to account for the kinetic energy decay, and further non-linear shape evolution, rather than assuming a stationary flow. Kinetic energy decay has recently been studied numerically [25], and power laws close to the predicted $t^{-10/7}$ have been observed in Ref. [40]. There is also predicted to be an associated $t^{2/7}$ evolution of the integral scale, which was not included in Ref. [25]. Other simulations (see Ref. [40]) show that the compressional mode spectrum evolves towards a self-similar form, which can be fitted by a broken power law with asymptotic indices $k^{5}$ at low wave numbers and $k^{-1}$ at high wave numbers. The evolution towards this shape also needs to be investigated, as the majority of the gravitational wave emission will probably occur during the evolutionary phase.
@@
Our results for the decorrelation functions also have implications for future modelling. In current implementations, it is assumed that the decorrelation functions of compressional modes oscillate indefinitely with period set by sound speed, leading to a kernel which (in the limit $kR_{*}\to\infty$) can be approximated by a $\delta$-function. The new form given in Eq. 65 improves the prediction of the growth rate of gravitational wave power spectra. The ratio between a $M_{\parallel}=1.14$ and $M_{\parallel}$ for the $C[U_{\parallel},U_{\parallel}]$ is contained in the range $[0.4,2.0]$, where the lowest ratios occur for the lowest values of $kR_{*}$ (below order 10), while the largest are closer to $kR^{*}\approx 9\times 10^{2}$.
@@
$\Omega_{\text{gw},0}/(H_{\text{n}}R_{*})^{2}\\ =\left\{\begin{array}[]{cc}(4.8\pm 1.1)\times 10^{-8}&(v_{\text{w}}=0.92,\alpha_{\text{n}}=0.67)\\ (1.3\pm 0.2)\times 10^{-8}&(v_{\text{w}}=0.44,\alpha_{\text{n}}=0.50)\\ \end{array}\right.$. (80)
@@
We conclude with a remark on the deflagration. As previously mentioned, due to reheating effects in front of the bubble wall, the wall speed is reduced from $v_{\text{w}}\approx 0.4\rightarrow$ to $v_{\text{w}}\approx 0.1$ in the course of the simulation. Despite this reducing the observed gravitational wave power, as explained in Ref. [37], this may provide a way to enact baryogenesis in the latter part of the phase transition. Electroweak baryogenesis [56] in its original form was argued to be most efficient in transitions with low asymptotic wall speeds, at odds with strong gravitational wave production. However, this reheating implies that electroweak baryogenesis and gravitational wave production need not be mutually exclusive: the two different regimes (prior and after reheating) could allow for a sufficiently strong (observable) gravitational wave imprint, while still allowing for baryogenesis to occur. For example, the deflagration we simulate produces a loud signal for sufficiently large $H_{\text{n}}R_{*}$, while approximately 30% of the universe is swept by walls with speed $v_{\text{w}}\approx 0.1$, which if maintained from the beginning would produce a much weaker signal. A detailed exploration of this effect as function of transition strength, wall speed and bubble size deserves further investigation.